Sales $ 35,000,000 Operating expenses Variable expenses Fixed expenses $28,000,000 3,500,000 Total expenses 31,500,000 Operating profit $ 3,500,000 1. Determine the breakeven point in sales dollars Breakeven point in sales dollars 17,500,000 2. Determine the required sales in dollars to earn a before-tax profit of $4,500,000. (Do not round intermediate calculations. Round your answer to the nearest whole dollar amount.) Required sales in dollars $ 40,000,000 3. What is the breakeven point in sales dollars if the variable cost increases by 10%? (Do not round intermediate calculations. Round your answer to the nearest whole dollar amount.) Breakeven point in sales dollars
Answer #1

Contribution margin=Sales-Variable expenses

=(35,000,000-28,000,000)=$7,000,000

Contribution margin ratio=Contribution margin/Sales

=(7,000,000/35,000,000)=20%

a.Breakeven point=Fixed expenses/Contribution margin ratio

=(3,500,000/0.2)=$17,500,000

b.Required Contribution margin=Fixed expenses+Target profit

=(3,500,000+4,500,000)=$8,000,000

Hence required sales=(8,000,000/0.2)=$40,000,000

c.New variable cost=(28,000,000*1.1)=$30,800,000

Contribution margin=Sales-Variable expenses

=(35,000,000-30,800,000)=$4,200,000

Hence Contribution margin ratio=(4,200,000/35,000,000)=12%

Hence breakeven sales=(3,500,000/0.12)

=$29,166,667(Approx).